OpenStudy (anonymous):

Please help? Suppose h:x→x2 −x.Then,h:−4→____ ? . A –20 C 12 B –8 D 20

jimthompson5910 (jim_thompson5910):

h: x --> x^2 - x is the same as saying h(x) = x^2 - x

jimthompson5910 (jim_thompson5910):

when you say h: -4 --> ___ you basically want to find h(-4)

OpenStudy (anonymous):

So (4x^2)-(-4)?

OpenStudy (anonymous):

*(-4x^2)-(-4)

jimthompson5910 (jim_thompson5910):

h(x) = x^2 - x h(-4) = (-4)^2 - (-4) h(-4) = 16 + 4 h(-4) = 20

jimthompson5910 (jim_thompson5910):

So -4 maps to 20 which means h: −4→__20__

OpenStudy (anonymous):

Ohh okay.

jimthompson5910 (jim_thompson5910):

in goes x = -4 out comes h(x) = 20

OpenStudy (anonymous):

Thanks :)

jimthompson5910 (jim_thompson5910):

yw